# Interpolating Video-LLMs: Toward Longer-sequence LMMs in a Training-free Manner

### v1

@@ Abstract
Advancements in Large Language Models (LLMs) inspire various strategies for integrating video modalities. A key approach is Video-LLMs, which incorporate an optimizable interface linking sophisticated video encoders to LLMs. However, due to computation and data limitations, these Video-LLMs are typically pre-trained to process only short videos, limiting their broader application for understanding longer video content. Additionally, fine-tuning Video-LLMs to handle longer videos is cost-prohibitive. Consequently, it becomes essential to explore the interpolation of Video-LLMs under a completely training-free setting. In this paper, we first identify the primary challenges in interpolating Video-LLMs: the video encoder and modality alignment projector are fixed, preventing the integration of additional frames into Video-LLMs, and the LLM backbone is limited in its content length capabilities, which complicates the processing of an increased number of video tokens. To address these challenges, we propose a specific INTerPolation method for Video-LLMs (INTP-Video-LLMs). We introduce an alternative video token rearrangement technique that circumvents limitations imposed by the fixed video encoder and alignment projector. Furthermore, we introduce a training-free LLM context window extension method to enable Video-LLMs to understand a correspondingly increased number of visual tokens. We analyze the deployment costs of INTP-Video-LLM, and find its efficiency bottleneck is on its KV cache cost. Accordingly, we introduce a training-free KV-cache compression mechanism that reduces memory overhead during inference. INTP-VideoLLM not only supports the processing of longer video sequences but also optimizes memory usage during inference—all achieved without the need for additional training. In practice, whereas pre-trained Video-LLaVA [Lin et al., 2023] models are configured to process just 8 frames, INTP allows these models to comprehend 32 frames.
@@ 1 Introduction
Large Language Models (LLMs) [OpenAI, 2023, Touvron et al., 2023a; b], have shown surprising capabilities in question answering and reasoning. Building on this foundation, Vision LLMs extend these abilities to include images, employing a vision encoder and an LLM to generate text responses given an image and a related question [Liu et al., 2023a, Zhang et al., 2024]. Recent advancements aim to extend this capability from whole-image understanding to video understanding [Lin et al., 2023, Zhang et al., 2023, Kim et al., 2024]. Video-LLMs combine video data with language models by integrating learnable interfaces that capture both spatial and temporal representative of video. Typically, these interfaces use a projection network [Li et al., 2023b, Maaz et al., 2023, Li et al., 2023a] to transform video content into video tokens that can be interpreted by LLMs, thereby bridging the gap between video information and text processing capabilities of LLMs (see Fig.1).
@@
However, existing Video-LLMs [Li et al., 2023b, Lin et al., 2023, Maaz et al., 2023, Li et al., 2023a, Kim et al., 2024] struggle to process long videos. In particular, Video-LLMs lack the temporal resolution necessary to interpolate precision time, limiting its application to long video [Huang et al., 2024]. For instance, Video-LLaMA [Li et al., 2023b] and Video-LLaVA [Lin et al., 2023] only sample 8 frames uniformly across an entire video, which often falls short for detailed temporal analysis [Huang et al., 2024]. Simply feeding more frames into these models does not resolve the issue, as their architectures cannot effectively process a larger number of frames. Consequently, there is an urgent need for Video-LLMs that can understand larger numbers of video frames.
@@
To reach this goal, one could consider training Video-LLMs from scratch to handle more extensive video frames. Unfortunately, this approach is impractical due to prohibitive training costs and data accessibility issues. On the one hand, the Video-LLMs consume more and more data, which making the accessibility of data more and more difficult. In some cases, the training data even are not released. Sometimes, the weights of those Video-LLM are released while the data is not open. On the other hand, the computation cost of training a Video-LLM is expensive. For example, even the relatively less costly Video-LLaVA [Lin et al., 2023] requires thousands of hours on Nvidia A100 GPUs. Therefore, retraining a Video-LLM to process more frames is not feasible in many scenarios.
@@
Figure 1: (Left) Video-LLMs consist of three main components: a video encoder, an alignment projector layer, and a fine-tuned LLM backbone. The process begins with the video encoder transforming video frames into a series of visual tokens. A projector then maps these tokens, aligning visual features with text features. The resulting aligned features, along with text prompts, are subsequently fed into the LLM backbone for visual understanding. (Right) INTP, a training-free Video-LLM interpolation technique, addresses the constraints of Video-LLMs. We employ an alternative video token rearrangement that bypasses the limitations set by the fixed video encoder and alignment projector. Additionally, we implement a training-free LLM context window interpolation method to allow Video-LLMs to process an increased number of visual tokens effectively.
@@
In this paper, we aim to enable Video-LLMs to understand videos with an extensive number of frames in a training-free manner. We first delineate the primary challenges associated with interpolating Video-LLMs: the video encoder and modality alignment projector are fixed, which prevents the integration of additional frames, and the content length capacity of the LLM backbone is limited, hindering the processing of an increased number of visual tokens.
@@
To overcome these obstacles, we propose a specific interpolation method for Video-LLMs, dubbed INTP. Addressing the challenge, we propose an alternative video token rearrangement technique that bypasses the restrictions imposed by the fixed video encoder and alignment projector. This approach allows us to utilize the pre-trained video encoder and alignment projector to generate an unlimited number of video tokens, while maintain the video tokens’ consistency. For the challenge, derived from the positional embedding mechanism in Video-LLMs (i.e., Rotary Position Embedding, RoPE [Su et al., 2024]), we develop a training-free Video-LLM context window extension method. This design ensures that the interpolated Video-LLM can handle any number of video frames.
@@
Upon developing the INTP-Video-LLM, we examined its deployment constraints and identified that the handling of an increased number of video tokens incurs additional memory usage. This finding echos the memory bottleneck in the long-sequence LLM deployment [Yuan et al., 2023; 2024]. To address this issue and enhance the deployment efficiency of INTP-Video-LLM, we introduce a training-free KV-cache compression technique that reduces memory overhead during inference. Consequently, the INTP-Video-LLM not only facilitates the processing of longer video sequences but also optimizes memory usage during inference, all without necessitating further training.
@@ 2 Related Work
Video Language Models. Prior to the emergence of LLMs, Yang et al. [2022] introduce FrozenBiLM, which combined a frozen vision encoder with a bidirectional language model for efficient video processing. In the LLM era, video-language models continue to evolve. For instance, Video-ChatGPT enhances video instruction tuning using high-quality instructional data [Maaz et al., 2023]. Li et al. [2023a] propose VideoChat, in which they utilize cross-attention mechanisms to integrate video tokens with user queries and dialogue context. Then, Li et al. [2024] propose the follow-up work, VideoChat2, in which they advance this integration with a multi-stage bootstrapping method, focusing on modality alignment and further instruction tuning. Video-LLaVA [Lin et al., 2023] introduces a pre-aligned encoder that supports both image and video modalities, enabling shared projections and joint training across tasks. However, handling long videos is very challenging due to the high computational complexity and memory demands of representing extensive video content with video tokens. To address these issues, various advanced temporal modeling techniques have been deployed. Chat-UniVi [Jin et al., 2023] proposes a unified model that dynamically merges spatial and temporal tokens using k-NN techniques to streamline processing. LLaMA-VID [Li et al., 2023b] utilize a dual token system to effectively compress video tokens. Despite these explorations, all previous methods have relied on some form of training. In contrast, our work is the first to explore the extension of the Video-LLM context window in a training-free manner, offering an approach to handling long video sequences without the computational expense and data requirements of traditional training methods.
@@
Long-context LLMs. LLMs are generally pre-trained within a specified context length, with models like LLaMA and LLaMA2 employing 2k and 4k tokens respectively [Touvron et al., 2023a; b]. Training LLMs from scratch to handle extended contexts is often prohibitively expensive for many researchers. Nevertheless, recent innovations have focused on extending the context length through fine-tuning methods. Recent advancements focus on modifying the position embedding (PE), particularly RoPE [Su et al., 2024], employed in LLMs such as LLaMA and Mistral [Jiang et al., 2023]. One of the main strategies is embedding scaling [Chen et al., 2023, Liu et al., 2023b, Peng et al., 2023]. Embedding scaling techniques modify position embeddings to accommodate longer sequences, ensuring they remain within the pre-training scope and preventing feature extrapolation. For instance, Chen et al. [2023] condense position indices to maintain alignment with the pre-training range, effectively expanding LLaMA’s context to 16,000 tokens with minimal fine-tuning, requiring only 1,000 steps. In a different approach, Liu et al. [2023b], Roziere et al. [2023] adjust the rotary base of RoPE, known as “NTK-aware” scaling. To the best of our knowledge, our work is the first exploration to extend the context window of Video-LLMs in a training-free manner.
@@
Post-training Compression for KV Cache. Recently, the compression of Key-Value (KV) caches has garnered significant attention due to the prohibitively high memory consumption associated with generating long contextual sequences in LLMs. Current methods can be briefly categorized into three types: Quantization-aware compression, eviction-based compression and attention-based compression. Quantization-aware compression reduces KV cache size by substituting the original KV cache with lower-precision approximations. KVQuant [Hooper et al., 2024], KIVI [Liu et al., 2024] and WKVQuant [Yue et al., 2024] are pioneering studies in KV cache quantization, revealing that keys and values should be quantized along different dimensions. Notably, KIVI compresses the KV cache to as few as 2 bits. Yang et al. [2024], Dong et al. [2024] further enhance quantization performance by identifying and selectively preserving more significant keys and values. SKVQ [Duanmu et al., 2024] advances upon KIVI using a sliding window technique, wherein quantization parameters are determined on a window-wise basis. ZipCache [He et al., 2024] establish a robust baseline for KV cache quantization, achieving compression ratios as high as 5x with negligible performance degradation. In this wrok, we analyse the development bottlenecks of long-context Video-LLMs, and propose using the quantization techniques to address the computational overhead challenges.
@@ 3 Method: Interpolating Video-LLMs
In this section, we first review the basic implementation of Video Large Language Models (Video-LLMs), with a particular focus on the main components and their training pipelines. We highlight the difficulties of obtaining a long-video-LLM from scratch (Sec.3.1). Next, we present a totally training-free method specifically designed for Video-LLM INTerPolation, called INTP. There are three advancements in the development and optimization of INTP: (1) A video encoder and modality-alignment projector extension method (Sec.3.2); (2) A video-LLM context window extension method (Sec. 3.3); and (3) An inference bottleneck analysis and a post-training quantization solution (Sec.3.4).
3.1 Preliminaries: Video Large Language Models (Video-LLM)
Video-LLMs are typically composed of three core modules: a video encoder, an alignment projector layer, and a fine-tuned large language model (LLM) backbone. The process begins with a video encoder—often a Vision Transformer (ViT)—which converts video input $\mathbf{X}_{v}\in\mathbb{R}^{N\times W\times H}$ into a series of visual tokens $\mathbf{Z}_{v}$, in which $N$ is number of frames, $W$ and $H$ is the pixel width and height of each frame. These tokens are processed by the input projector $\mathbf{\Theta}_{\mathbf{Z}_{v}\rightarrow\mathbf{H}_{v}}$, which maps the encoded visual features to the text feature space $\mathbf{H}$. The aligned features, along with text prompts $\mathbf{H}_{q}$, are then input into a LLM backbone, and then the language model can generate the response. The architecture of a typical Video-LLMs is illustrated in Fig. 1.
Video encoders are crucial for compressing the raw video into compact representations. These encoders are often trained from scratch via multimodal alignment pretraining [Yin et al., 2023]. For instance, LanguageBind [Zhu et al., 2023] employs contrastive learning during the pretraining phase to gradually align video modality with language modality. Given that LLMs inherently process text, bridging the modality gap between natural language and video is essential. However, training a Video-LLM end-to-end is prohibitively expensive. A more feasible approach involves using a learnable connector between the pre-trained visual encoder and the LLM, along with fine-tuning the LLM to better interpret visual tokens. Although those components can be trained via instruction tuning [Gong et al., 2023] or alignment tuning [Sun et al., 2023], their training remains costly. For example, to training the video encoder, Zhu et al. [2023] collect a large-scale video-and-text dataset, VIDAL-10M. In addition, even without pre-training the video encoder, fine-tuning the Video-LLaVA [Lin et al., 2023] still consumes approximately 200 hours on Nvidia A100 GPUs.
@@ 3.1 Preliminaries: Video Large Language Models (Video-LLM)
In practice, the challenges in extending the Video-LLMs can be summarized from two key aspects: the fixed nature of the video encoder and modality alignment projector, and the limited content length capacity of the LLM backbone. Firstly, the video encoder and the alignment projector are typically configured during initial training and remain unchanged thereafter. This prevents the system from incorporating additional frames beyond the preset limit, which limits the model’s ability to adapt to videos of different lengths or to capture more granular temporal details within prolonged video sequences. Secondly, the LLM backbone, which is designed to process textual information, has a predefined limit on the number of tokens it can handle. This constrains the model’s capacity to interpret a larger number of visual tokens, effectively capping the amount of visual information that can be processed. As the complexity and length of videos increase, these constraints become significant bottlenecks, hindering the model’s ability to fully understand and generate coherent responses based on longer video inputs.

### v2

@@ Abstract
Advances in Large Language Models (LLMs) have inspired various strategies for integrating video modalities. A key approach is Video-LLMs, which incorporate an optimizable interface linking sophisticated video encoders to LLMs. However, due to computation and data limitations, existing Video-LLMs are typically pre-trained to process only short videos, limiting their broader application for understanding longer video content. Additionally, fine-tuning Video-LLMs to handle longer videos is cost-prohibitive. Consequently, it is essential to explore the interpolation of Video-LLMs under a completely training-free setting. In this paper, we first identify the primary challenges in interpolating Video-LLMs: the video encoder and modality alignment projector are fixed, preventing the integration of additional frames into Video-LLMs, and the LLM backbone is limited in its content length capabilities, which complicates the processing of an increased number of video tokens. To address these challenges, we propose an INTerPolation method for Video-LLMs (INTP-Video-LLMs). We introduce a video token rearrangement technique that circumvents limitations imposed by the fixed video encoder and alignment projector. Furthermore, we introduce a training-free LLM context window extension method to enable Video-LLMs to understand a correspondingly increased number of visual tokens. We analyze the deployment costs of INTP-Video-LLM, and find its efficiency bottleneck is on its KV cache cost. Accordingly, we introduce a training-free KV-cache compression mechanism that reduces memory overhead during inference. INTP-VideoLLM not only supports the processing of longer video sequences but also optimizes memory usage during inference—all achieved without the need for additional training. In practice, whereas pre-trained Video-LLaVA [Lin et al., 2024] models are configured to process just 8 frames, INTP allows these models to comprehend 32 frames.
@@ 1 Introduction
Large Language Models (LLMs) [OpenAI, 2023, Touvron et al., 2023a; b], have shown tremendous capabilities in question answering and reasoning. Building on this foundation, Vision LLMs extend these abilities to include images, employing a vision encoder and an LLM to generate text responses given an image and a related question [Liu et al., 2023a, Zhang et al., 2024]. Recent advancements aim to extend this capability from image understanding to video understanding [Lin et al., 2024, Zhang et al., 2023a, Kim et al., 2024]. Video-LLMs combine video data with language models by integrating learnable interfaces that capture both spatial and temporal information in video. Typically, these interfaces use a projection network [Li et al., 2023b, Maaz et al., 2023, Li et al., 2023a] to transform video content into video tokens that can be interpreted by LLMs, thereby bridging the gap between video information and text processing capabilities of LLMs (see Fig.1).
@@
However, existing Video-LLMs [Li et al., 2023b, Lin et al., 2024, Maaz et al., 2023, Li et al., 2023a, Kim et al., 2024] struggle to process long videos. In particular, Video-LLMs often lack the temporal resolution necessary to precisely model temporal events, limiting its application to long video [Huang et al., 2024]. For instance, Video-LLaMA [Li et al., 2023b] and Video-LLaVA [Lin et al., 2024] only sample 8 frames uniformly across an entire video, which is often too short for detailed temporal analysis [Huang et al., 2024]. Simply feeding more frames into these models does not resolve the issue, as their architectures cannot effectively process a larger number of frames. Consequently, there is a pressing need for Video-LLMs that can understand larger numbers of video frames.
@@
To reach this goal, one could consider training Video-LLMs from scratch to handle more extensive video frames. Unfortunately, this approach can be impractical due to prohibitive training costs and data accessibility issues. On the one hand, Video-LLMs require increasingly large datasets, making data acquisition more challenging, especially considering copyright restrictions on video-language paired data. In some cases, the training data are not publicly available. Sometimes, only the model weights are released while the data is not open. On the other hand, the computation cost of training a Video-LLM is expensive. For example, even the relatively less costly Video-LLaVA [Lin et al., 2024] requires thousands of hours on Nvidia A100 GPUs. Therefore, retraining a Video-LLM to process more frames is not feasible in many scenarios.
@@
Figure 1: (Left) Video-LLMs consist of three main components: a video encoder, an alignment projector layer, and a fine-tuned LLM backbone. The process begins with the video encoder transforming video frames into a series of visual tokens. A projector then maps these tokens, aligning visual features with text features. The resulting aligned features, along with text prompts, are subsequently fed into the LLM backbone for visual understanding. (Right) INTP, a training-free Video-LLM interpolation technique, addresses the existing constraints of Video-LLMs. We employ a video token rearrangement that bypasses the limitations set by the fixed video encoder and alignment projector. Additionally, we implement a training-free LLM context window interpolation method to allow Video-LLMs to process an increased number of visual tokens effectively.
@@
In this paper, we aim to significantly increase the number of frames that existing Video-LLMs can use to understand videos in a training-free manner. We first delineate the primary challenges associated with interpolating Video-LLMs: the video encoder and modality alignment projector are fixed, which prevents the integration of additional frames, and the content length capacity of the LLM backbone is limited, hindering the processing of an increased number of visual tokens.
@@
To overcome these obstacles, we propose an interpolation method for Video-LLMs, dubbed INTP. For the challenge, we propose a video token rearrangement technique that bypasses the restrictions imposed by the fixed video encoder and alignment projector. This approach allows utilizing the pre-trained video encoder and alignment projector to generate an increased number of video tokens for LLM reasoning while maintaining the video tokens’ temporal consistency. For the challenge, derived from the positional embedding mechanism in Video-LLMs (i.e., Rotary Position Embedding, RoPE [Su et al., 2024]), we develop a training-free Video-LLM context window extension method. This design ensures that the interpolated Video-LLM can handle any number of video frames.
@@
Upon developing the INTP-Video-LLM, we examine its deployment constraints and identify that the handling of an increased number of video tokens incurs additional memory usage. This finding echos the memory bottleneck in long-sequence LLM deployment [Yuan et al., 2023; 2024]. To address this issue and enhance the deployment efficiency of INTP-Video-LLM, we introduce a training-free KV-cache compression technique that reduces memory overhead during inference. Consequently, INTP-Video-LLM not only facilitates the processing of longer video sequences but also optimizes memory usage during inference, all without necessitating further training.
@@ 2.1 Video Language Models
Prior to the emergence of LLMs, Yang et al. [2022] introduced FrozenBiLM, which combined a frozen vision encoder with a bidirectional language model for efficient video processing. In the LLM era, video-language models continue to evolve. For instance, Video-ChatGPT enhances video instruction tuning using high-quality instructional data [Maaz et al., 2023]. Li et al. [2023a] propose VideoChat, which utilizes cross-attention mechanisms to integrate video tokens with user queries and dialogue context. Li et al. [2024] propose a follow-up work, VideoChat2, which advances this integration with a multi-stage bootstrapping method, focusing on modality alignment and further instruction tuning. Video-LLaVA [Lin et al., 2024] introduces a pre-aligned encoder that supports both image and video modalities, enabling shared projections and joint training across tasks. However, handling long videos is very challenging due to the high computational complexity and memory demands of representing extensive video content with video tokens. To address these issues, various advanced temporal modeling techniques have been deployed. Chat-UniVi [Jin et al., 2023] proposes a unified model that dynamically merges spatial and temporal tokens using k-NN techniques to streamline processing. LLaMA-VID [Li et al., 2023b] uses a dual token system to effectively compress video tokens. Despite these explorations, all previous methods rely on some form of training. In contrast, our work is the first to explore the extension of the Video-LLM context window in a training-free manner, offering an approach to handling long video sequences without the computational expense and data requirements of traditional training methods.
2.2 Long-context LLMs
@@
LLMs are generally pre-trained with a specified context length, with models like LLaMA and LLaMA2 employing 2k and 4k tokens respectively [Touvron et al., 2023a; b]. Training LLMs from scratch to handle extended context is often prohibitively expensive for many researchers. Thus, recent innovations have focused on extending the context length through fine-tuning methods. Recent advancements focus on modifying the position embedding (PE), particularly RoPE [Su et al., 2024], employed in LLMs such as LLaMA and Mistral [Jiang et al., 2023]. One of the main strategies is embedding scaling [Chen et al., 2023, Liu et al., 2023b, Peng et al., 2023], which modifies position embeddings to accommodate longer sequences, ensuring they remain within the pre-training scope and preventing feature extrapolation. For instance, Chen et al. [2023] condense position indices to maintain alignment with the pre-training range, effectively expanding LLaMA’s context to 16,000 tokens with minimal fine-tuning, requiring only 1,000 steps. In a different approach, Liu et al. [2023b], Roziere et al. [2023] adjust the rotary base of RoPE, known as “NTK-aware” scaling. To the best of our knowledge, our work is the first exploration to extend the context window of Video-LLMs in a training-free manner.
2.3 Post-training Compression for KV Cache
@@
Recently, the compression of Key-Value (KV) caches has garnered significant attention due to the prohibitively high memory consumption associated with generating long contextual sequences in LLMs. Current methods can be briefly categorized into three types: Quantization-aware, eviction-based, and attention-based. Quantization-aware compression reduces KV cache size by substituting the original KV cache with lower-precision approximations. KVQuant [Hooper et al., 2024], KIVI [Liu et al., 2024] and WKVQuant [Yue et al., 2024] are pioneering studies in KV cache quantization, revealing that keys and values should be quantized along different dimensions. Notably, KIVI compresses the KV cache to as few as 2 bits. Yang et al. [2024], Dong et al. [2024] further enhance quantization performance by identifying and selectively preserving more significant keys and values. SKVQ [Duanmu et al., 2024] advances upon KIVI using a sliding window technique, wherein quantization parameters are determined on a window-wise basis. ZipCache [He et al., 2024] establishes a robust baseline for KV cache quantization, achieving compression ratios as high as 5x with negligible performance degradation. In this work, we analyze the development bottlenecks of long-context Video-LLMs and identify that the major constraint lies in the LLM’s KV cache. To address this computational overhead, we leverage quantization techniques, effectively optimizing the model’s memory usage and inference efficiency.
@@ 3 Method: Interpolating Video-LLMs
In this section, we first review the standard implementation of Video Large Language Models (Video-LLMs), with a particular focus on the main components and training pipelines. We highlight the difficulties of obtaining a long-video-LLM from scratch (Sec.3.1). Next, we present a totally training-free method specifically designed for Video-LLM INTerPolation, called INTP. There are three advancements in the development and optimization of INTP: (1) A video encoder and modality-alignment projector extension method (Sec.3.2); (2) A video-LLM context window extension method (Sec. 3.3); and (3) An inference bottleneck analysis and a post-training quantization solution (Sec.3.4).
3.1 Preliminaries: Video Large Language Models (Video-LLMs)
Video-LLMs are typically composed of three core modules: a video encoder, an alignment projector layer, and a fine-tuned large language model (LLM) backbone. The process begins with a video encoder—often a Vision Transformer (ViT)—which converts video input $\mathbf{X}_{v}\in\mathbb{R}^{N\times W\times H}$ into a series of visual tokens $\mathbf{Z}_{v}$, in which $N$ is number of frames, $W$ and $H$ is the pixel width and height of each frame. These tokens are processed by the input projector $\mathbf{\Theta}_{\mathbf{Z}_{v}\rightarrow\mathbf{H}_{v}}$, which maps the encoded visual features to the text feature space $\mathbf{H}$. The aligned features, along with text prompts $\mathbf{H}_{q}$, are then input into a LLM backbone, and then the language model can generate the response. The architecture of a typical Video-LLMs is illustrated in Fig. 1 (left).
Video encoders are crucial for compressing the raw video into compact representations. These encoders are often trained from scratch via multimodal alignment pretraining [Yin et al., 2023]. For instance, LanguageBind [Zhu et al., 2023] employs contrastive learning during the pretraining phase to gradually align video modality with language modality. Given that LLMs inherently process text, bridging the modality gap between natural language and video is essential. However, training a Video-LLM from scratch is prohibitively expensive. A more feasible approach involves using a learnable connector between the pre-trained visual encoder and the LLM, along with fine-tuning the LLM to better interpret visual tokens. Although those components can be trained via instruction tuning [Gong et al., 2023] or alignment tuning [Sun et al., 2023], their training remains costly. For example, to train the video encoder, Zhu et al. [2023] collect a large-scale video-and-text dataset, VIDAL-10M. In addition, even without pre-training the video encoder, fine-tuning the Video-LLaVA [Lin et al., 2024] still consumes approximately 200 hours on Nvidia A100 GPUs.
@@ 3.1 Preliminaries: Video Large Language Models (Video-LLMs)
In practice, the challenges in extending the Video-LLMs can be summarized from two key aspects: the fixed nature of the video encoder and modality alignment projector, and the limited content length capacity of the LLM backbone. First, the video encoder and alignment projector are typically configured during initial training and remain unchanged thereafter. This prevents the system from incorporating additional frames beyond the preset limit, which limits the model’s ability to adapt to videos of different lengths or to capture more granular temporal details within prolonged video sequences. Second, the LLM backbone has a predefined limit on the number of tokens it can handle. This constrains the model’s capacity to interpret a larger number of visual tokens, effectively capping the amount of visual information that can be processed. As the complexity and length of videos increase, these constraints become significant bottlenecks, hindering the model’s ability to fully understand and generate coherent responses based on longer video inputs.